.The probability that a patient recovers from a delicate heart operation is 0.8. What is the probability that: a) exactly 2 of the next 3 patients who have this operation will survive? b) all of the next 3 patients who have this operation survive?


1
Expert's answer
2021-06-16T14:28:46-0400

Binomial distribution

p=0.8

n=3

"P(x)=\binom{n}{x}P^x(1-P)^{n-x}, x=0,1,2...,n."

"a. P(x=2)"

"P(x=2)=\binom{3}{2}0.8^20.2=0.384"

b. "P(x=3)"

"P(x=3)=\binom{3}{3}0.8^3=0.512"